OpenStudy (anonymous):

simplify (v/w^-2)^-3

OpenStudy (anonymous):

\[\left(\frac{v}{w^{-2}}\right)^{-3}=\frac{(v)^{-3}}{(w^{-2})^{-3}}=\frac{v^{-3}}{w^{(-2\times-3)}}=\frac{v^{-3}}{w^6}\]